OpenStudy (anonymous):

please help and explain: solve for x: 4/x-3+5/x+3=6x^2-9

OpenStudy (tkhunny):

You'll have to try again. Remember the convention, "Order of Operations". Do you mean \(\dfrac{6}{x^{2} - 9}\) on the right-hand side? Your task is to add the fractions. Find the Least Common Denominator and proceed!

OpenStudy (anonymous):

i still dont get it

OpenStudy (tkhunny):

You must first answer the clarifying question. What did you mean on the right hand side? Practice. Please add these: 1/3 + 1/2 = ??? Demonstrate your methods.

OpenStudy (anonymous):

2.5

OpenStudy (tkhunny):

2.5 what? Is that the answer to the right-hand side? Is that the METHOD used to add the fractions? Please make an effort to communicate better. 1/3 + 1/2 is nowhere near 2.5. Please demonstrate your METHOD to obtain your answer. What did you do first? What did you do second? What was next...?

OpenStudy (anonymous):

my bad 2/5

OpenStudy (tkhunny):

Also incorrect. Until you show your METHOD, no one can help you. Additional Practice: Add these -- 1 Cat + 1 Dog = ??? Seriously. It's not a joke.

OpenStudy (anonymous):

2

OpenStudy (tkhunny):

You added a cat and a dog and managed to get the number 2? Does that make any sense to you at all? Where did the dog and cat go?

OpenStudy (anonymous):

ok whatever my grandma was first my mom was second and I am third

OpenStudy (tkhunny):

In other words, you're just not serious about learning this part of mathematics. Why did you bother to type in the problem?

OpenStudy (anonymous):

so you would ask

OpenStudy (anonymous):

I just dont understand can you please explain it little bit better sorry

OpenStudy (tkhunny):

You must answer my question. Have you EVER successfully added two fractions with unlike denominators? Add: 1 dog + 1 cat = ?? The correct answer is: "That's silly. You can't add dogs and cats." Add: 1/3 + 1/2 = ?? A good first answer might be, "That's silly. How can you add thirds to halves" A good second answer might be, "I wonder of I can change them so that the denominators would be the same? Then I COULD add them. Numbers aren't the same as household pets. With numbers, you have a little latitude." A good third answer might be: "The least common multiple of 2 and 3 is 6. Let's change the problem to 6ths. Add: 2/6 + 3/6 = 5/6. Done!"

OpenStudy (anonymous):

oh ok